There are 9 counters in a bag. 7 of the counters are green. 2 of the counters are blue. Two counters are chosen at random, what is the probability one counter of each colour is chosen.

There are two methods to pick one counter of each. Either pick the (1)green counter first and then blue, or a (2)blue counter first and then green. We must calculate the probability of both of these scenarios and then add the two together for the answer.(1) Probability of picking Green and then Blue: 7/9 x 2/8 = 14/72(In the second probability the denominator is 8 as we have already picked a counter making the total number of counters in the bag 8)(2) Probability of picking Blue and then Green:2/9 x 7/8 = 14/72(Once again the same rule applies as above. The second denominator is one lower as a counter has been removed from the bag)To find out the probability of picking one of each colour, we must add the probabilities of the two scenarios together:14/72 + 14/72 = 28/72Answer = 28/72

UB

Related Maths GCSE answers

All answers ▸

I struggle with the following type of question: "The first four terms of an arithmetic sequence are 5, 9, 13, 17. Write down an expression, in terms of n, for the nth term in the sequence." How should I approach this?


P is a point on the circle with equation x^2 + y^2 = 80. P has x-coordinate 4 and is below the x-axis.Work out the equation of the tangent to the circle at P.


Solve the Simultaneous equations x^2 + y^2 =29, y-x=3.


Solve the simultaneous equation: 3x+y =-4 and 3x-4y=6